OpenStudy (anonymous):

A man took 1/4 th of a pie which was 3 times the part of pie the other family members had. Then how many family members are there?

OpenStudy (anonymous):

There should be 11 family members because if taking 1/4 of the pie is 3 times the size that the others took, that means that the others took 1/12 of the pie.

OpenStudy (anonymous):

what is the equation ?

OpenStudy (anonymous):

I made a mistake there should be 10.

OpenStudy (anonymous):

How did you get 10 ?

OpenStudy (anonymous):

Because 1/4 is 3 times the size of 1/12, right?

OpenStudy (anonymous):

@skullpatrol ...can you tell me the equation.....just the setting up part

OpenStudy (anonymous):

so if everyone were taking same sized slices (1/12) there would be twelve people, but 1 person took 3 times that size, so its like taking three slices... then everyone finished the pie taking 1/12 sized slices... that means there are 10 people

OpenStudy (anonymous):

am I wrong?

OpenStudy (anonymous):

Don't ask me....I am still trying to figure it out.

OpenStudy (anonymous):

@ajprincess ... @ash2326 ...

OpenStudy (anonymous):

@Jonask ....

OpenStudy (anonymous):

Do you know how o set up the equation ?

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@Mertsj ..

OpenStudy (anonymous):

I should know this....my brain is just not fully awake yet..lol

OpenStudy (anonymous):

You don't always need an equation to solve these type of questions...

OpenStudy (anonymous):

I guess your right....

OpenStudy (anonymous):

yeah!!...the answer is 10..gr8 work

OpenStudy (mertsj):

The rest of the family ate 3/4 pie. And 1/4 = 3x so they each ate 1/12 pie. 1/12(# of family members)=3/4 # of family members = 9 9+ the person who ate 1/4 of the pie = 10 family members.

OpenStudy (anonymous):

thank you ..